1 / 18


PS. Dept meeting 1/12/17. People -. Two new faces – D. Tapp – Linac G. Donaldson - Booster People should have goal settings entered in workday Merit increase have gone out – I think. Pre acc update. Source A: Still having issues with high extractor spark rates

Download Presentation


An Image/Link below is provided (as is) to download presentation Download Policy: Content on the Website is provided to you AS IS for your information and personal use and may not be sold / licensed / shared on other websites without getting consent from its author. Content is provided to you AS IS for your information and personal use only. Download presentation by click this link. While downloading, if for some reason you are not able to download a presentation, the publisher may have deleted the file from their server. During download, if you can't get a presentation, the file might be deleted by the publisher.


Presentation Transcript

  1. PS Dept meeting 1/12/17

  2. People - • Two new faces – • D. Tapp – Linac • G. Donaldson - Booster • People should have goal settings entered in workday • Merit increase have gone out – I think

  3. Pre acc update • Source A: • Still having issues with high extractor spark rates • Had failed extractor cable • Currently having solenoid gas valve problems • High leak rate • Large excursions in pressure • Rebuilt valve, pulser seems OK at this time • Source B: • Operational since startup • Some internal sparking (anode-cathode) • Beam current low 60mA because of extraction gap • Failed extractor cable • Projects: • Current regulated arc modulator: • Working in test stand, final testing before sending boards out to get made • Solenoid gas valve pulser: • Initial design running in test stand. • A couple of mods need to be checked out then have boards made • Need to order supplies (0-300V) • Designing circuit to detect extractor sparks, then inhibit DTI trigger • Should greatly reduce the amount of energy dumped into the source during sparks

  4. Linac Report Jan/2016 Jan/2017 Linac Overview Linac Report Department Meeting

  5. Linac Report Linac Overview <EffLinac> =~ 90% <Ibeam> = 23.2 mA <Uptime> = 97.5% Linac Report Department Meeting

  6. - L3 Communications – San Carlo unit is closing down! Not profitable – stakeholders want to demolish the factory and sell the property – make more $$. • Decision was made to move the equipment to a branch in Pennsylvania. • Our vendor left L3 and moved to CPI. Situation is not very promising for long term support on this equipment. • 2017 Look-Ahead • PIP Work • Installation of Laser Notch system • Installation of 2 Marx Modulator • Proton Source • Maintenance on the LE and HE stations Genera l Ac t i v i t ies Highlights Good news… • Installation of full size Marx Modulator in LRF5 • Marx working flawlessly in LRF1 And not so good news… • KRF2 pulse transformer (PT#1) • Rebuilt by Stangenes ($$$) (twice) • Klystron • S/N 004 rebuilt by L3 • Yet to be tested S/N 004 & PT#1 combination does not hold voltage!(< 90kV) • S/N 017 – failed (gun assembly) • S/N 05-R1 being rebuilt • S/N 016 – next to be rebuilt Last year news Linac Report Department Meeting 1/12/17

  7. Overview of last 30 days of Booster running HEP

  8. Booster Beam to Numi 11/11/16 – 01/11/2017

  9. Booster Beam to BNB • 11/11/16 – 01/11/2017

  10. Booster Downtime 12/11/2016 – 01/11/2017 • Excellent uptime! Averaging < 2 hours of downtime per week. • Running more beam intensity out of Booster at a faster rep rate, ~13Hz, then ever before. • Maintaining a beam efficiency > 90%

  11. Total Protons Delivered by Booster over one week, 12/11 – 01/11

  12. Booster BLM plot while running beam to Numi and BNB

  13. What’s Happening in Booster? • New blood in Booster Group, welcome Geoff Donaldson Operations: • Always tuning to reduce losses • Improving cogging with MI/RI • Making changes to allow laser notching starting next month • Running on a new digital damper system with new software • Commissioning new LLRF system • Installing BRF21 BRF cavity, currently testing for 20Hz • Building cavity 22, install during long shutdown or before (hoping) Studies: • Measuring tunes (LOCO) to improve the Booster orbit for much efficiency (coming soon) • One-Shot Bhat studies to improve capture and transition crossing • Collimator studies to optimize/redesign collimator system to reduce losses. Current studies plan being written.

  14. FY17 • PIP continues • Goal of 2.4E17 pph • New shielding assessment • Start to operate at ~14 Hz • Laser notcher • Power switch to master sub • Feb 15,16 • Summer shutdown planning starting – Joe! • Two Modulators • Two Booster cavities • Beyond – PIP 1+ • Depending upon funding • More flux out of PS • Can we reach 2.7E17 pph • 800 KW on NOvA target? • Can we reach 900 KW? • What direction will the DOE go? • PIP II funding? • LBNE funding? • Lab Funding?

  15. ES&H - Safety • We are now one lab – safety completely re-arranged. • AD really does not have a ES&H department anymore – reports to 7th floor safety • New faces – most of whom I do not know • PS has not had any notable safety issues • Safety walk through with new people in about one month – will see how it goes….

  16. Budget - • Really up in the air – • I have no idea what is going to happen • Signs are not promising • Will do what I can to keep people informed

  17. Mech Support • MS – rearranged their department • Did not help us • We have a new ME department engineer and water tech for Booster • We had no real input into this – just told who… • I want to know about MS requests/issues for our department

  18. Misc • 50th anniversary • May require us to improve Linac/pre-acc gallery

More Related